OpenStudy (anonymous):

is the number 72 a prime or composite number

OpenStudy (idealist10):

72 is not a prime number. 36*2=72 8*9=72 12*6=72 1*72=72 18*4=72

OpenStudy (anonymous):

composite number.

OpenStudy (anonymous):

composite, not a prime.

OpenStudy (kainui):

Any number that ends in an even number is even, and if a number is even, you can divide it by 2. Similarly, any number when you add up its digits adds up to 3, 6 or 9 is divisible by 3.

OpenStudy (bibby):

Prove it @7Aeslehc7 @markiaa15 You can't just say something and hope the statement is self evident. What is a composite number? More importantly, WHY isn't 72 prime?

OpenStudy (anonymous):

it is composite composite are not prime

OpenStudy (anonymous):

A prime number is a whole number that only has two factors which are itself and one. A composite number has factors in addition to one and itself. The numbers 0 and 1 are neither prime nor composite. All even numbers are divisible by two and so all even numbers greater than two are composite numbers.

OpenStudy (anonymous):

72 is a composite number because 72 has divisors rather than 1 and itself. 72 can be divided by 2 and 2.

OpenStudy (anonymous):

did that help?

OpenStudy (bibby):

brava

OpenStudy (anonymous):

you welcome.